OpenStudy (anonymous):

What is the square root of 49 over the square root of 2?

hero (hero):

Hint: \[\sqrt{49} = 7\]

OpenStudy (anonymous):

so is it 7 over the square root of 2?

hero (hero):

Before rationalization, yes

hero (hero):

After rationalization: \[\frac{7\sqrt{2}}{2}\]

hero (hero):

Do you know how to rationalize @mscrosscountry ?

OpenStudy (mertsj):

\[\frac{7}{\sqrt{2}}\times\frac{\sqrt{2}}{\sqrt{2}}=\frac{7\sqrt{2}}{2}\]